I am having problems with this question. can u please help me?
determine the odds for and odds against each event for the given probability: P(H)= 1/4
thank u very much!

Answer by sudhanshu_kmr(1152)

P(H) = 1/4
P'(H) = 1 - P(H) = 1 -1/4 = 3/4
odds in favor for the event = P(H)/P'(H) = (1/4) /(3/4) = 1/3

odds against the event = P'(H)/P(H) = (3/4) /(1/4) = 3/1 = 3